# The JWST-NIRCam View of Sagittarius C. III. The Extinction Curve

### v1

@@ Abstract
Methods. We determined extinction ratios using two different methods, both based on measuring the reddening vector using the slope of red clump stars, whose intrinsic properties are well known, in observed colour-magnitude diagrams.
Results. The extinction curve derived in this work is in good agreement with previous results in the literature. We obtained the following extinction ratios relative to F162M: . Besides, we found different values of the extinction index for the short- (m, ) and long-wavelength (m, ) regimes, with the extinction curve flattening at longer wavelengths. Comparison with extinction curves derived both inside and outside the Galactic centre suggests that the infrared extinction curve does not significantly vary in the central regions, and shows no significant evidence for variations between different lines of sight beyond the inner Galaxy within the uncertainties.
@@
The Galactic centre (GC) represents a unique astrophysical laboratory, as it is the closest galactic nucleus and the only one where individual stars can be resolved down to milliparsec scales (e.g. Genzel et al., 2010). However, studying its stellar content and structure is challenging due to the extreme stellar crowding and high interstellar extinction, which can exceed mag (e.g. Scoville et al., 2003; Nishiyama et al., 2008; Fritz et al., 2011; Schödel et al., 2010). This severe extinction restricts observations to the infrared regime, where high-angular resolution photometry is essential to mitigate crowding effects (e.g. Nogueras-Lara et al., 2018). Although intrinsic stellar colours in the near-infrared are relatively small (e.g. Schödel et al., 2010; Nogueras-Lara et al., 2018), their variation across spectral types complicates photometric classification (e.g. Schödel et al., 2014).
@@
An additional difficulty is that different wavelength dependencies of the extinction have been reported in the literature (e.g Nishiyama et al., 2006; Schödel et al., 2010; Fritz et al., 2011; Nogueras-Lara et al., 2019a). It was generally assumed that the extinction curve in the m range follows a power-law of the form (e.g. Nishiyama et al., 2008; Fritz et al., 2011), where is the wavelength and is the extinction index. Previous studies found that (e.g. Nishiyama et al., 2006; Schödel et al., 2010; Alonso-García et al., 2017; Nogueras-Lara et al., 2018), although more recent work (e.g. Hosek et al., 2018; Nogueras-Lara et al., 2019a, 2020) shows that the extinction index varies with wavelength. One of the most detailed spectroscopic studies of the extinction curve towards the GC was presented by Fritz et al. (2011), who used hydrogen emission lines from the Short Wave Spectrometer of the Infrared Space Observatory (ISO-SWS; de Graauw et al., 1996) and spectroscopy from SINFONI, an integral-field spectrograph mounted on the VLT (Eisenhauer et al., 2003). While they found an approximate power-law in the m, the extinction curve longwards of m appears to be complex (e.g. Lutz, 1999).
@@
An accurate determination of the extinction curve requires studying the spectral energy distribution (SED) of stars with known properties. Ideally, this should be done using narrow band filters and across a broad wavelength range to avoid non-linear photometric effects associated with broad-band filters (for further details see Jones & Hyland, 1980; Straižys & Lazauskaitė, 2008; Maíz Apellániz et al., 2020; Nogueras-Lara et al., 2021b). This makes the Near-Infrared Camera (NIRCam; Rieke et al. 2005) onboard the James Webb Space Telescope (JWST) an ideal instrument for extinction studies as it combines high angular resolution with multi-filter coverage from 1 to 5 m.
@@
In this context, the GC Sagittarius C (Sgr C) region was recently observed with NIRCam (Crowe et al., 2025; Bally et al., 2025). Sgr C is a star-forming region including ionized (HII) regions located on the western edge of the Milky Way’s nuclear stellar disc (e.g. Lang et al., 2010; Kendrew et al., 2013). Embedded in a dense molecular cloud, it lies in a region of high extinction ( mag; Nogueras-Lara, 2024), making it an excellent target for deriving the extinction curve (e.g. Henshaw et al., 2023). In this paper, we use the NIRCam observations of Sgr C to derive the extinction curve in a wavelength range from 1.15 to 4.8 m.
@@ 2 Observations and data reduction
Observations of Sgr C were conducted with JWST-NIRCam on September 2023 (Program ID: 4147, PI: S. Crowe). NIRCam consists of two redundant modules (A and B) separated by a central gap of , each providing a field of view of and containing both a short wavelength (SW, –m) and a long wavelength (LW, –m) channel (Burriesci, 2005). Each module contains four SW detectors with inter-detector gaps of – and a pixel scale of /pixel, along with two LW detectors with a pixel scale of /pixel (Burriesci, 2005). To ensure full spatial coverage and mitigate detector gaps, six dither positions were employed using the FULLBOX 6TIGHT dither pattern, resulting in a final effective field of view of approximately . The observations used four SW filters (F115W, F162M, F182M, and F212N) and four LW filters (F360M, F405N, F470N, and F480M).
The data were reduced using the JWST standard pipeline (jwst v1.12.5; Bushouse et al. 2023), which includes detector corrections (Stage 1), calibrations (Stage 2), and image combination (Stage 3). Figure 1 shows a false-colour RGB image of Sgr C created from Stage 3 mosaics. Residual noise, not fully addressed by the pipeline, was corrected using image1overf.py111https://github.com/chriswillott/jwst/blob/master/image1overf.py, following the method described in Crowe et al. (2025). Throughout this work, we use Stage 2 images with noise removed, because the photometric zero point of each NIRCam detector is different. We use the corresponding Vega magnitude zero points provided by the STScI222NRC_ZPs_1126pmap.txt, to be found in https://jwst-docs.stsci.edu/jwst-near-infrared-camera/nircam-performance/nircam-absolute-flux-calibration-and-zeropoints.. For further details on the filter set, observational setup and data reduction of the Sgr C NIRCam data, we refer to Crowe et al. (2025).
@@ 3.1 Point spread function fitting photometry
We carried out point spread function (PSF) fitting photometry on the Stage 2 NIRCam images, i.e., before combining them into mosaics. This approach was chosen because the photometric zero points vary significantly between different NIRCam detectors. This also allowed us to estimate robust uncertainties from independent measurements because the majority of sources were observed multiple times, that is, in images with different dither offsets and even with different detectors.
@@
We extracted the image, its corresponding error map, and the pixel map with the data quality flags from the multi-extension fits files. The error map is automatically generated by the JWST Stage 2 pipeline and contains the uncertainty estimates for each pixel, provided as standard deviations. These uncertainties are computed by combining the variance due to Poisson noise and read noise, which are also included as separate extensions in the FITS file. We created saturation masks for each image by using the corresponding pixel data quality flags. A bad pixel map was created from the non-saturated pixels that were flagged as hot, cold, or ”DO_NOT_USE”. We filled all bad and saturated pixels in the error map with the median of the ten highest valid values of the error map, which was necessary for subsequent processing. The precise value used to fill the error map did not have any significant effect on the outcome.
@@
Subsequently, we used the stpsf tool to create theoretical PSFs for each detector and filter (https://stpsf.readthedocs.io/en/latest/). Saturated stars were repaired by fitting the theoretical PSFs to their unsaturated wings, after having subtracted bright, but unsaturated stars from their vicinity. For this, we used a modified version of the repair_saturated routine from the StarFinder software package (Diolaiti et al., 2000). Repairing the saturated stars proved to be essential to minimise the photometric contamination of nearby stars and to avoid the detection of spurious sources in the diffraction spikes of the saturated stars.
@@
Finally, we used StarFinder to carry out PSF fitting photometry and astrometry, because it is optimised for crowded fields. We modified StarFinder to always assume positive extended emission throughout the image, because this makes physical sense for NIRCam in this extremely bright field. The correlation threshold was set to 0.85, the detection threshold to 3 with two iterations, the background fitting box to 11 pixels and the deblend parameter was set to 1. A correlation threshold of 0.85 is stricter than the standard 0.7 of StarFinder, but well justified given the very well defined and stable PSF of NIRCam (see, e.g., Schödel, 2010; Gallego-Cano et al., 2018). After tests varying these parameters, we found that they provided the optimum trade-off between high completeness, avoidance of detecting spurious sources, and speed in the source detection process. As concerns the background box, it can be varied without any significant impact on the results. The previously mentioned error map was used as an uncertainty map for the procedure. Finally, we used the NIRCam zero points of each filter to convert from calibrated fluxes to magnitudes.
@@ 3.2 Final catalogues
We built the final catalogues using the function match_to_catalog_sky from the Python package Astropy (version 5.2.2, Astropy Collaboration et al., 2013, 2018, 2022) to identify common stars in the overlap regions between the detectors. For each common star, we computed the mean value of parameters such as its magnitude and coordinates. We chose a match radius of , which is smaller than the diffraction limit for all filters, but large enough to accommodate also fainter stars that show greater astrometric scatter (see Sect. 2). Given that systematic astrometric offsets between chips are below , the expected astrometric uncertainties are significantly smaller than our matching threshold. Therefore, a search radius of ensures robust cross-matching without introducing spurious matches.
@@ 3.3 Photometric uncertainties
In the final catalogues, some stars are detected only once, while others have multiple detections. In the latter case, for each detection, we first derived the flux from the computed magnitude. Then, we estimated the flux uncertainty, following a similar approach to that of Nogueras-Lara et al. (2018):
@@
We used the uncertainties provided by the StarFinder package, which fits the stars with an IDL implementation of the Newton-Gauss method and uses the provided pixel uncertainty map to formally estimate the uncertainties. Stars with lower signal-to-noise ratio thus have larger uncertainties (Fig. 2 top). Since this procedure depends on providing an accurate pixel uncertainty map, we cross checked the uncertainties by using the measurements of stars with multiple detections. As the bottom panel of Fig. 2 shows, the photometric uncertainties from multiply detected stars can be very small due to the random nature of this method (e.g. when two measurements are very similar), but the upper envelope of the cloud of uncertainties is in good agreement with the uncertainties provide by StarFinder, thus demonstrating the validity of the uncertainties delivered by StarFinder.
@@ 3.4 Colour-magnitude diagrams
To build the colour-magnitude diagrams (CMDs), we cross-matched the catalogues from selected pairs of filters using a distance of . Figure 3 shows the F480M vs. F162M-F480M CMD. The purple box indicates the red clump (RC) feature, identified as an overdensity in the diagram. It consists of giant stars burning helium in their cores (Girardi, 2016). Due to their weak dependence on age and metallicity, the intrinsic properties of these stars, which are well characterised, make them ideal for tracing extinction and determining distances. Moreover, since the RC is a prominent feature in the CMD, we can see clearly how it extends parallel to the reddening vector (black arrow in Fig. 3). The latter was computed using the extinction curve derived in this work (see Sect. 4.1).
We also observe the presence of a bright and a faint RC. This double RC feature has been interpreted as evidence for two stellar populations formed in separate star formation episodes—an older component (8 Gyr) and a younger one (1 Gyr) (see Nogueras-Lara et al., 2019a). Additionally, we consider stars with F162M-F480M mag to belong to the foreground population (e.g. Nogueras-Lara et al., 2021a).
@@ 4.1 Reference filter method
The slope of the RC features can be used to derive extinction ratios, as it corresponds to (Nishiyama et al., 2006), where is the extinction in magnitudes at a given wavelength , and is the colour excess, defined as
From this relation, the extinction ratio can be expressed as
where is the slope of the RC in the vs. CMD. Thus, for the reference filter method, we adopt . To derive the slope of the RC features, we applied the methodology described in Nogueras-Lara et al. (2019b) and Nogueras-Lara et al. (2020). Namely, we divided the RC region into 5 to 7 narrow vertical bins, depending on the colour width of the RC features. We show the selection boxes and binning used for each CMD in the reference filter method in Fig. 8. The selection boxes were defined to ensure a relatively homogeneous density of RC stars across the CMDs, minimising the impact of incompleteness at the faint end. We then verified the presence of a double RC sequence (see Sect. 3.4) by analysing the magnitude distribution within each bin, fitting both a single-Gaussian and a two-Gaussian model using the GaussianMixture (GMM) function from the SCIKIT-LEARN Python package (Pedregosa et al., 2011). In all cases, we found that the two-Gaussian model describes best the underlying distribution, as determined by the Bayesian information criterion (Schwarz, 1978) and the Akaike information criterion (Akaike, 1974).
For each bin, we calculated the representative colour as the median of the stellar colours within the bin, and estimated its uncertainty as the standard error of the median (). The peaks of the Gaussians identified by the GMM provide the y-axis positions of the RC features—that is, the magnitude values corresponding to the bright and faint RC populations in each bin. The uncertainty was computed as the standard error of the mean of each Gaussian distribution. The resulting points and their uncertainties are shown in Fig. 3 for the F480M vs. F162M-F480M CMD as an example. The Gaussian fit results for the remaining CMDs are shown in Appendix B.
@@
The number of bins (and consequently their width): we tested using two fewer bins than in the nominal case (e.g. 4 if it was 6), as well as 10 bins in all cases.
@@
The test on cross-match radius also accounts for the different angular resolutions between the short- and long-wavelength channels, though it had a negligible effect on the results. Variations in the RC box along the colour axis further helped to evaluate completeness effects: larger boxes include the less dense—and therefore less complete—edges of the RC, while smaller boxes focus on its densest, most complete region. Table 1 shows the individual contributions to the systematic uncertainty for each filter and for both the bright and faint RC features. The total systematic uncertainty was computed as the quadratic sum of all components. The variation of the RC selection box yields the largest contribution to the total systematic uncertainty, as shown in Table 1.
@@
In addition, we interpolated the extinction curve derived with the reference filter method, along with its associated uncertainties, using the CubicSpline function from the SciPy library (Virtanen et al., 2020). For each filter, we derived the effective wavelength, , as described in Appendix A. As part of this process, we also computed the extinction index between each filter and the reference filter. The resulting values are shown in Tables 2 and 3 for both the bright and faint RC, along with their average.

### v2

@@ Abstract
Methods. We determined extinction ratios using two different methods, both based on measuring the reddening vector using the slope of red clump stars (whose intrinsic properties are well known) in observed colour-magnitude diagrams.
Results. The extinction curve derived in this work is in good agreement with previous results in the literature. We obtained the following extinction ratios relative to F162M: . Additionally, we found different values of the extinction index for the short- (m, ) and long-wavelength (m, ) regimes, with the extinction curve flattening at longer wavelengths. Comparison with extinction curves derived both inside and outside the Galactic centre suggests that the infrared extinction curve does not significantly vary in the central regions, and shows no significant evidence of variations between different lines of sight beyond the inner Galaxy within the uncertainties.
@@
The Galactic centre (GC) represents a unique astrophysical laboratory, as it is the closest galactic nucleus and the only one where individual stars can be resolved down to milliparsec scales (e.g. Genzel2010). However, studying its stellar content and structure is challenging due to the extreme stellar crowding and high interstellar extinction, which can exceed mag (e.g. Scoville2003; Nishiyama2008; Fritz2011; Schoedel2010). This severe extinction restricts observations to the infrared regime, where high-angular-resolution photometry is essential to mitigate crowding effects (e.g. NoguerasLara2018). Although intrinsic stellar colour values in the near-infrared are relatively small (e.g. Schoedel2010; NoguerasLara2018), their variation across spectral types complicates photometric classification (e.g. Schoedel2014).
@@
An additional difficulty is that different wavelength dependences of the extinction have been reported in the literature (e.g Nishiyama2006; Schoedel2010; Fritz2011; NoguerasLara2019a). It was generally assumed that the extinction curve in the m range follows a power law of the form (e.g. Nishiyama2008; Fritz2011), where is the wavelength and is the extinction index. Previous studies found that (e.g. Nishiyama2006; Schoedel2010; AlonsoGarcia2017; NoguerasLara2018), although more recent work shows that the extinction index varies with wavelength (e.g. Hosek2018; NoguerasLara2019a; NoguerasLara2020). One of the most detailed spectroscopic studies of the extinction curve towards the GC was presented by Fritz2011, who used hydrogen emission lines from the Short Wave Spectrometer of the Infrared Space Observatory (ISO-SWS; Graauw1996) and spectroscopy from the Spectrograph for INtegral Field Observations in the Near Infrared (SINFONI; Eisenhauer2003), an integral-field spectrograph mounted on the Very Large Telescope (VLT). While they found an approximate power law in the m, the extinction curve longwards of m appears to be complex (e.g. Lutz1999).
@@
An accurate determination of the extinction curve requires studying the spectral energy distribution of stars with known properties. Ideally, this should be done using narrow-band filters and across a broad wavelength range to avoid non-linear photometric effects associated with broad-band filters (for further details see Jones1980; Straizys2008; MaizApellaniz2020; NoguerasLara2021a). This makes the Near-Infrared Camera (NIRCam; Rieke2005 Rieke2005) on board the James Webb Space Telescope (JWST) an ideal instrument for extinction studies as it combines high angular resolution with multi-filter coverage from 1 to 5 m.
@@
In this context, the GC Sagittarius C (Sgr C) region was recently observed with NIRCam (Crowe2025; Bally2025). Sgr C is a star-forming region that hosts ionised (HII) regions; it is located on the western edge of the Milky Way’s nuclear stellar disc (e.g. Lang2010; Kendrew2013). Embedded in a dense molecular cloud, it lies in a region of high extinction ( mag; NoguerasLara2024), making it an excellent target for deriving the extinction curve (e.g. Henshaw2023). In this study we used the NIRCam observations of Sgr C to derive the extinction curve in the wavelength range 1.15 to 4.8 m.
@@ 2 Observations and data reduction
Observations of Sgr C were conducted with JWST-NIRCam on 22 September 2023 (programme ID: 4147, PI: S. Crowe). NIRCam consists of two redundant modules (A and B) separated by a central gap of , each providing a field of view of and containing both a short-wavelength (SW; –m) and a long-wavelength (LW; –m) channel (Burriesci2005). Each module contains four SW detectors with inter-detector gaps of – and a pixel scale of /pixel, along with two LW detectors with a pixel scale of /pixel (Burriesci2005). To ensure full spatial coverage and mitigate detector gaps, six dither positions were employed using the FULLBOX 6TIGHT dither pattern, resulting in a final effective field of view of approximately . The observations used four SW filters (F115W, F162M, F182M, and F212N) and four LW filters (F360M, F405N, F470N, and F480M).
The data were reduced using the JWST standard pipeline (jwst v1.12.5; Bushouse2023 Bushouse2023), which includes detector corrections (Stage 1), calibrations (Stage 2), and image combination (Stage 3). Figure 1 shows a false-colour red-green-blue image of Sgr C created from Stage 3 mosaics. Residual noise, not fully addressed by the pipeline, was corrected using image1overf.py111https://github.com/chriswillott/jwst/blob/master/image1overf.py, following the method described in Crowe2025. Throughout this work we used Stage 2 images with noise removed, because the photometric zero point of each NIRCam detector is different. We used the corresponding Vega magnitude zero points provided by the STScI222NRC_ZPs_1126pmap.txt, to be found in https://jwst-docs.stsci.edu/jwst-near-infrared-camera/nircam-performance/nircam-absolute-flux-calibration-and-zeropoints.. For further details on the filter set, observational setup and data reduction of the Sgr C NIRCam data, we refer to Crowe2025.
@@ 3.1 Point spread function fitting photometry
We carried out point spread function (PSF) fitting photometry on the Stage 2 NIRCam images, i.e. before combining them into mosaics. This approach was chosen because the photometric zero points vary significantly between different NIRCam detectors. This also allowed us to estimate robust uncertainties from independent measurements because the majority of sources were observed multiple times, that is, in images with different dither offsets and even with different detectors.
@@
We extracted the image, its corresponding error map, and the pixel map with the data quality flags from the multi-extension fits files. The error map is automatically generated by the JWST Stage 2 pipeline and contains the uncertainty estimates for each pixel, provided as standard deviations. These uncertainties are computed by combining the variance due to Poisson noise and read noise, which are also included as separate extensions in the FITS file. We created saturation masks for each image by using the corresponding pixel data quality flags. A bad pixel map was created from the non-saturated pixels that were flagged as hot, cold, or ‘DO_NOT_USE’. We filled all bad and saturated pixels in the error map with the median of the ten highest valid values of the error map, which was necessary for subsequent processing. The precise value used to fill the error map did not have any significant effect on the outcome.
@@
Subsequently, we used the stpsf333https://stpsf.readthedocs.io/en/latest/ tool to create theoretical PSFs for each detector and filter. Saturated stars were repaired by fitting the theoretical PSFs to their unsaturated wings, after having subtracted bright, but unsaturated stars from their vicinity. For this, we used a modified version of the repair_saturated routine from the StarFinder software package (Diolaiti2000). Repairing the saturated stars proved to be essential to minimise the photometric contamination of nearby stars and to avoid the detection of spurious sources in the diffraction spikes of the saturated stars.
@@
Finally, we used StarFinder to carry out PSF fitting photometry and astrometry, because it is optimised for crowded fields. We modified StarFinder to always assume positive extended emission throughout the image, because this makes physical sense for NIRCam in this extremely bright field. The correlation threshold was set to 0.85, the detection threshold to 3 with two iterations, the background fitting box to 11 pixels and the deblend parameter was set to 1. A correlation threshold of 0.85 is stricter than the standard 0.7 of StarFinder, but well justified given the very well-defined and stable PSF of NIRCam (see e.g. Schoedel2010a; GallegoCano2018). After tests varying these parameters, we found that they provided the optimum trade-off between high completeness, avoidance of detecting spurious sources, and speed in the source detection process. As concerns the background box, it can be varied without any significant impact on the results. The previously mentioned error map was used as an uncertainty map for the procedure. Finally, we used the NIRCam zero points of each filter to convert from calibrated fluxes to magnitudes.
@@ 3.2 Final catalogues
We built the final catalogues using the function match_to_catalog_sky from the Python package Astropy (version 5.2.2; AstropyCollaboration2013; AstropyCollaboration2018; AstropyCollaboration2022) to identify common stars in the overlap regions between the detectors. For each common star, we computed the mean value of parameters such as its magnitude and coordinates. We chose a match radius of , which is smaller than the diffraction limit for all filters, but large enough to accommodate also fainter stars that show greater astrometric scatter (see Sect. 2). Given that systematic astrometric offsets between chips are below , the expected astrometric uncertainties are significantly smaller than our matching threshold. Therefore, a search radius of ensures robust cross-matching without introducing spurious matches.
@@ 3.3 Photometric uncertainties
In the final catalogues, some stars are detected only once, while others have multiple detections. In the latter case, for each detection, we first derived the flux from the computed magnitude. Then, we estimated the flux uncertainty, following a similar approach to that of NoguerasLara2018:
@@
We used the uncertainties provided by the StarFinder package, which fits the stars with an IDL implementation of the Newton-Gauss method and uses the provided pixel uncertainty map to formally estimate the uncertainties. Stars with lower signal-to-noise ratios thus have larger uncertainties (Fig. 2 top). Since this procedure depends on providing an accurate pixel uncertainty map, we cross-checked the uncertainties by using the measurements of stars with multiple detections. As the bottom panel of Fig. 2 shows, the photometric uncertainties from multiply detected stars can be very small due to the random nature of this method (e.g. when two measurements are very similar), but the upper envelope of the cloud of uncertainties is in good agreement with the uncertainties provide by StarFinder, thus demonstrating the validity of the uncertainties delivered by StarFinder.
@@ 3.4 Colour-magnitude diagrams
To build the colour-magnitude diagrams (CMDs), we cross-matched the catalogues from selected pairs of filters using a distance of . Figure 3 shows the F480M versus F162M-F480M CMD. The purple box indicates the red clump (RC) feature, identified as an overdensity in the diagram. It consists of giant stars burning helium in their cores (Girardi2016). Due to their weak dependence on age and metallicity, the intrinsic properties of these stars, which are well characterised, make them ideal for tracing extinction and determining distances. Moreover, since the RC is a prominent feature in the CMD, we can see clearly how it extends parallel to the reddening vector (black arrow in Fig. 3). This vector was computed using the extinction curve derived in this work (see Sect. 4.1).
We also observe the presence of a bright and a faint RC. This double RC feature has been interpreted as evidence of two stellar populations formed in separate star formation episodes—an older component (8 Gyr) and a younger one (1 Gyr; see NoguerasLara2019a). Additionally, we considered stars with F162M-F480M mag to belong to the foreground population (e.g. NoguerasLara2021).
@@ 4.1 Reference filter method
The slope of the RC features can be used to derive extinction ratios, as it corresponds to (Nishiyama2006), where is the extinction in magnitudes at a given wavelength , and is the colour excess, defined as . From this relation, the extinction ratio can be expressed as
where is the slope of the RC in the versus CMD. Thus, for the reference filter method, we adopted . To derive the slope of the RC features, we applied the methodology described in NoguerasLara2019 and NoguerasLara2020. Namely, we divided the RC region into five to seven narrow vertical bins, depending on the colour width of the RC features. We show the selection boxes and binning used for each CMD in the reference filter method in Fig. 8. The selection boxes were defined to ensure a relatively homogeneous density of RC stars across the CMDs, minimising the impact of incompleteness at the faint end. We then verified the presence of a double RC sequence (see Sect. 3.4) by analysing the magnitude distribution within each bin, fitting both a single-Gaussian and a two-Gaussian model using the GaussianMixture (GMM) function from the SCIKIT-LEARN Python package (Pedregosa2011). In all cases, we found that the two-Gaussian model describes best the underlying distribution, as determined by the Bayesian information criterion (Schwarz1978) and the Akaike information criterion (Akaike1974).
For each bin, we calculated the representative colour as the median of the stellar colours within the bin, and estimated its uncertainty as the standard error of the median (). The peaks of the Gaussians identified by the GMM provide the y-axis positions of the RC features—that is, the magnitude values corresponding to the bright and faint RC populations in each bin. The uncertainty was computed as the standard error of the mean of each Gaussian distribution. The resulting points and their uncertainties are shown in Fig. 3 for the F480M versus F162M-F480M CMD as an example. The Gaussian fit results for the remaining CMDs are shown in Appendix B.
@@
The number of bins (and consequently their width): we tested using two fewer bins than in the nominal case (e.g. using four if it was six), as well as ten bins in all cases.
@@
The test on the cross-match radius also accounts for the different angular resolutions between the SW and LW channels; however, this test had a negligible effect on the results. Variations in the RC box along the colour axis further helped us evaluate completeness effects: larger boxes include the less dense—and therefore less complete—edges of the RC, while smaller boxes focus on its densest, most complete region. Table 1 shows the individual contributions to the systematic uncertainty for each filter and for both the bright and faint RC features. The total systematic uncertainty was computed as the quadratic sum of all components. The variation of the RC selection box yields the largest contribution to the total systematic uncertainty, as shown in Table 1.
@@
In addition, we interpolated the extinction curve derived with the reference filter method, along with its associated uncertainties, using the CubicSpline function from the SciPy library (Virtanen2020). For each filter, we derived the effective wavelength, , as described in Appendix A. As part of this process, we also computed the extinction index between each filter and the reference filter. The resulting values are shown in Tables 2 and 3 for both the bright and faint RC, along with their average.